Let x denote the time taken to run a road race. Suppose x is approximately normally distributed with a mean of 190 minutes and a standard deviation of 21 minutes. If one runner is selected at random, what is the probability that this runner will complete this road race in 221 to 240 minutes?

Round your answer to four decimal places.
Answer #1

Solution :

Given that ,

mean = \mu = 190

standard deviation = \sigma = 21

P(221 < x < 240) = P[(221 - 190)/ 21) < (x - \mu ) /\sigma  < (240 - 190) / 21) ]

= P(1.48 < z < 2.38)

= P(z < 2.38) - P(z < 1.48)

Using z table,

= 0.9913 - 0.9306

= 0.0607